OpenStudy (anonymous):

Which expression is equivalent to (7^3)−2? 1 over 7 times 7 times 7 times 7 times 7 times 7 7 1 over 7 negative 1 over 7 times 7 times 7 times 7 times 7 times 7

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@IrishBoy123 @imqwerty

OpenStudy (anonymous):

I need help with some exponent questions if you can help that'd be awesome

OpenStudy (phi):

what does a negative exponent mean ? any idea ?

OpenStudy (anonymous):

Just like a regular exponent just negative?

OpenStudy (phi):

there is a long explanation of why it is so, but cutting to the chase \[ a^{-b}= \frac{1}{a^b} \] and \[ a^{b}= \frac{1}{a^{-b}} \]

OpenStudy (anonymous):

hmm i get confused with fractions and exponents so I really dont know

OpenStudy (phi):

that "rule" says: flip the fraction *and* negate the exponent.

OpenStudy (anonymous):

This is a combo of both lol

OpenStudy (anonymous):

hmm

OpenStudy (phi):

you can learn it if you have time try guessing what \( 2^{-1} \) is. (flip it and negative the exponent)

OpenStudy (anonymous):

its a negative exponent?

OpenStudy (phi):

\( 2^{-1} \) has an exponent of -1 you can "rewrite it" by 1) flipping it 2) make the exponent -(-1) = +1

OpenStudy (anonymous):

..

OpenStudy (anonymous):

the 1 is the exponent?

OpenStudy (anonymous):

right

OpenStudy (phi):

can you rewrite \( 2^{-1} \) ?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

maybe

OpenStudy (phi):

flip (invert) means if you have a, write 1/a if you have 1/a write a if you have (stuff) write 1/(stuff)

OpenStudy (anonymous):

so like 1-2?

OpenStudy (anonymous):

like the - sign is negative?

OpenStudy (anonymous):

i do not get it D:

OpenStudy (phi):

2^(-1) first step: FLIP 1/2^(-1)) second step: change the exponent -1 to -(-1) (which simplifies to 1) we get 1/2^1 or just 1/2

OpenStudy (phi):

\[ 2^{-1} = \frac{1}{2^1} = \frac{1}{2} \]

OpenStudy (anonymous):

ok so the 2^-1 would be 1^-2?

OpenStudy (anonymous):

or the 1/2

OpenStudy (anonymous):

ok i think i got that

OpenStudy (phi):

if you just had 3 \[ 3 \text{ flipped is } \frac{1}{3} \]

OpenStudy (anonymous):

so instead of 1^-3 itd be 1/3??

OpenStudy (phi):

ok, one more \[ 3^{-2} \] can you flip and negate the exponent?

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

2/3?

OpenStudy (anonymous):

or 2^-3?

OpenStudy (anonymous):

ya 2/3 that would be correct i believe

OpenStudy (phi):

none of those. you don't change 3^-2 when you flip it rather, we think of it as \[ \frac{3^{-2}}{1} \] and swap top and bottom to flip it. then change the -2 to +2

OpenStudy (anonymous):

mm

OpenStudy (anonymous):

its still really confusing

OpenStudy (phi):

let's try this flip 1/2

OpenStudy (anonymous):

1^2?

OpenStudy (anonymous):

1^-2

OpenStudy (anonymous):

im sorry D:

OpenStudy (phi):

in \( \frac{1}{2} \) what is the top number ?

OpenStudy (anonymous):

the 1

OpenStudy (phi):

and 2 is the bottom number. what do you get if you swap those ?

OpenStudy (anonymous):

2/1

OpenStudy (phi):

and what do you get if you flip 2/1 ?

OpenStudy (anonymous):

1/2?

OpenStudy (phi):

yes can you flip 1/3 ?

OpenStudy (anonymous):

ohhhh ok 31

OpenStudy (anonymous):

3/1*

OpenStudy (phi):

yes

OpenStudy (anonymous):

ahhhhh ok

OpenStudy (phi):

normally when we flip a number, say 2 to 1/2 we get different numbers. (2 is not 1/2) but if we *also* change its exponent, we get the same number

OpenStudy (anonymous):

mm

OpenStudy (phi):

in other words \[ 2^{-1}\] if we flip it , to get \( \frac{1}{2^{-1}} \) and then change the -1 to +1, \[ 2^{-1}=\frac{1}{2} \]

OpenStudy (anonymous):

ok so we pretty much eliminate the -1? if we do +1?

OpenStudy (phi):

another example \[ 2^{-5} = \frac{1}{2^5} \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then if we changed the exponent to +5 it'd go as 1/2?

OpenStudy (phi):

I don't understand the question. can you ask in a different way?

OpenStudy (anonymous):

hmm so the 1/2-5

OpenStudy (anonymous):

if we changed the -5 to +5 then instead of 1/2-5 it would be 1/2?

OpenStudy (phi):

\[ \frac{1}{2^{-5}} \] 1) flip it and 2) change the sign on the 5 we get \[ \frac{1}{2^{-5}} =2^5\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still really weird to me

OpenStudy (phi):

it all makes sense (once you learn what is going on) but for the moment, just learn the rules 2^5 can be written as 1/2^-5 2^-5 can be written as 1/2^5

OpenStudy (anonymous):

ok

OpenStudy (phi):

\[ (7^3)^{−2}\] can you rewrite this? (tread the (7^3) as one thing)

OpenStudy (phi):

*treat

OpenStudy (anonymous):

ok umm 3/7

OpenStudy (phi):

leave (7^3) alone. it is one thing. keep it one thing. but if we do (7^3)^ -2 what can we do to make the -2 positive ?

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

i believe flip the numbers around? I really dont know DD:

OpenStudy (phi):

yes flip. think of (7^3) as one "number"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 3/7^-2?

OpenStudy (anonymous):

or 3^7-2

OpenStudy (phi):

you changed (7^3) . don't change it.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

7^3 stays the same

OpenStudy (anonymous):

if we made it different it would be 7/3^2?

OpenStudy (phi):

yes, what changes is we write 1/(7^3)^2

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

ok so the 7 stays with the 3

OpenStudy (anonymous):

ok ok

OpenStudy (phi):

we now have \[ \frac{1}{(7^3)^2}\]

OpenStudy (anonymous):

ya

OpenStudy (phi):

do you know that x^2 means x*x ?

OpenStudy (anonymous):

no i didnt

OpenStudy (phi):

do you know that 3^2 means 3*3 ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

and 5^2 means 5*5

OpenStudy (anonymous):

yes

OpenStudy (phi):

what about (7^3)^2 any idea ?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

I was think 7x3x2 but that wouldnt be correct right